OpenStudy (anonymous):

What Is the answer to this question? 12÷4+13 ? 2+22÷2 Is it: A. < B. > C. = this is on my test so I need the answer quickly!! thx! Will fan and medal!

OpenStudy (anonymous):

b

OpenStudy (anonymous):

12/4+13= 3+13=16. 2+22/2=11+2=13. 16>13

OpenStudy (delaneyparris):

b